# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v4

@@ B.1 Formulating RMIA
Let $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ denote, respectively, the mean and standard deviation of $t_{b}=\ell(\mathcal{M}(b),y_{b})$ estimated from shadow models, where $a$ denotes which of $(\bm{x},y_{\bm{x}})$ and $(\bm{z},y_{\bm{z}})$ is in the training set. By Equation 11 in (Zarifzadeh et al., 2024), the pairwise likelihood ratio is given as
@@
$\mathrm{Score}_{\mathrm{RMIA}}(\bm{x})=\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(x,\bm{z})\geq\gamma\right),$ (A63)
@@ B.2 Per-example RMIA
$t_{\bm{x}}$ =$=\begin{cases}\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x},\bm{x}}Z&\text{ if }(x,y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{\bm{z},\bm{x}}+\sigma_{\bm{z},\bm{x}}Z&\text{ if }(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}\end{cases}$ (A67)
$t_{\bm{z}}$ =$=\begin{cases}\mu_{\bm{x},\bm{z}}+\sigma_{\bm{x},\bm{z}}Z&\text{ if }(x,y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}\\ \mu_{\bm{z},\bm{z}}+\sigma_{\bm{z},\bm{z}}Z&\text{ if }(x,y_{\bm{x}})\notin\mathcal{D}_{\mathrm{target}}\wedge(\bm{z},y_{\bm{z}})\in\mathcal{D}_{\mathrm{target}}.\end{cases}$ (A68)
@@
When $(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}$ and $(z,y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x}}Z$ and $t_{\bm{z}}=\mu_{\bm{x},\bm{z}}+\sigma_{\bm{z}}Z$. Thus, $\lambda$ becomes
@@
$\textsc{tpr}_{\mathrm{RMIA}}(x)$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{x}}\geq\log\gamma)\geq\tau\right)$ (A100)
@@
Note that, unlike per-example LiRA, we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations A74 and A75 depend on the parameters computed from shadow models.
@@ B.3 RMIA power-law
$\Psi(\bm{x},C)=\frac{\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle^{2}}{||\bm{x}||^{2}}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle^{2}}{||\bm{z}||^{2}}\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle^{2}}{||\bm{x}||^{2}}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle^{2}}{||\bm{z}||^{2}}\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}{\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{||\bm{x}||}+\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{||\bm{z}||}\right|\mid y_{\bm{z}}=y_{\bm{x}}\right]+(C-1)\mathbb{E}_{\bm{z}}\left[\left|\frac{\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle}{||\bm{x}||}+\frac{\langle\bm{z},\bm{z}-m_{\bm{z}}\rangle}{||\bm{z}||}\right|\mid y_{\bm{z}}\not=y_{\bm{x}}\right]}.$ (A108)
@@
To apply Lemma 5, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$. Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example. Using similar argument as in the proof of Theorem 2, we have
@@
$s_{y_{\bm{x}}}^{(\bm{z})}(\bm{x})$ =$\begin{cases}\frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}+\bm{z}-\bm{x}_{y_{\bm{x}},S}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \frac{1}{S}\langle\bm{x},\sum_{i=1}^{S}\bm{x}_{y_{\bm{x}},i}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A111)
@@
$\mu_{\bm{z},\bm{x}}$ =$\begin{cases}\langle\bm{x},m_{y_{\bm{x}}}\rangle+\frac{1}{S}\langle\bm{x},\bm{z}-m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}=y_{\bm{z}}\\ \langle\bm{x},m_{y_{\bm{x}}}\rangle\quad&\text{ if }y_{\bm{x}}\not=y_{\bm{z}}\end{cases}$ (A115)
@@
$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-z\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A125)
@@
Now that from Lemma 5 we have
@@
We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A142) is as tight as that for $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation A139) for sufficiently large $S$. Let us denote
@@
Hence $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})=O(1/\sqrt{S})$. On the other hand, from Equations A139 and A142 we have for sufficiently large $S$
@@
where $c_{1}$ is some constant. Since for large $S$, $\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]$ and $\tau$ are independent of $S$, Equation A156 scales as $O(1/\sqrt{S})$.
@@
As for the LiRA power-law, bounding $||\bm{x}-m_{\bm{x}}||$ and $||\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
@@ C.3 Datasets
Table A2 shows the datasets used in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $>80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 which is not part of the original VTAB benchmark.
@@ Appendix D Additional results
D.1 Additional results for Section 4
This Section contains additional results for Section 4.
@@ D.1.1 Vulnerability as a function of shots
This section displays additional results to Figure 1 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.1, A3 and A4.
[…]
@@ D.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure 2 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures A.2, A5 and A6.
[…]
@@ D.1.3 Data for FiLM and from scratch training
Table A7: MIA vulnerability data used in Figure 5(b). Note that the data from Carlini et al. (2022) is only partially tabular, thus we estimated the tpr at fpr from the plots in the Appendix of their paper.
@@ D.1.4 Predicting dataset vulnerability as function of $S$ and $C$
This section provides additional results for the model based on Equation 13
Table A8: Results for fitting Equation 13 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.952 0.907 -0.506 0.011 -44.936 $0$0.529 -0.484 0.01 0.946 0.854 -0.555 0.014 -39.788 $0$0.582 -0.527 0.001 0.930 0.790 -0.627 0.019 -32.722 $0$0.664 -0.589 $\beta_{C}$ (for $C$) 0.1 0.952 0.907 0.090 0.021 4.231 0.000 0.048 0.131 0.01 0.946 0.854 0.182 0.026 6.960 0.000 0.131 0.234 0.001 0.930 0.790 0.300 0.036 8.335 0.000 0.229 0.371 $\beta_{0}$ (intercept) 0.1 0.952 0.907 0.314 0.045 6.953 0.000 0.225 0.402 0.01 0.946 0.854 0.083 0.056 1.491 0.137 -0.027 0.193 0.001 0.930 0.790 -0.173 0.077 -2.261 0.025 -0.324 -0.022
Figure A.3 shows the performance for all considered fpr.
@@
[…]
D.2 Simpler variant of the prediction model
[…]
The general form this variant can be found in Equation A174, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
Table A9 provides tabular results on the performance of the variant.
Table A9: Results for fitting Equation A174 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
coeff. fpr $R^{2}$ test $R^{2}$ coeff. value std. error $t$ $p>|z|$ coeff. $[0.025$ coeff. $0.975]$ $\beta_{S}$ (for $S$) 0.1 0.908 0.764 -0.248 0.008 -30.976 $0$0.264 -0.233 0.01 0.940 0.761 -0.416 0.011 -36.706 $0$0.438 -0.393 0.001 0.931 0.782 -0.553 0.017 -32.507 $0$0.586 -0.519 $\beta_{C}$ (for $C$) 0.1 0.908 0.764 0.060 0.015 3.955 0.000 0.030 0.089 0.01 0.940 0.761 0.169 0.021 7.941 0.000 0.127 0.211 0.001 0.931 0.782 0.297 0.032 9.303 0.000 0.234 0.360 $\beta_{0}$ (intercept) 0.1 0.908 0.764 0.029 0.032 0.913 0.362 -0.034 0.093 0.01 0.940 0.761 -0.118 0.045 -2.613 0.010 -0.208 -0.029 0.001 0.931 0.782 -0.295 0.068 -4.345 $0$0.429 -0.161
Figure A.4 plots the performance of the variant similar to Figure 5(a) in the main text.
@@ D.2 Simpler variant of the prediction model
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation A174 fitted Table A3 (ViT-B). The triangles show the median tpr for the train set (ViT-B; Table A3) and circle the test set (R-50; Table A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
@@ D.3 Empirical results for RMIA
Figures A.5, A.6 and A.7 report additional results for RMIA Zarifzadeh et al. (2024).
[…]
@@
Table A10 displays the tabular results for Figure 6 in Section 4.4.
Table A10: Tabular results for Figure 6 on when attacking a ViT-B (Head) fine-tuned on PatchCamelyon. We display the median over six seeds at $\textsc{fpr}=0.1$.
@@
In Section 4.5, we compare the results of our empirical models of vulnerability (Sections 4.3 and 4.4) to DP bounds. Below we explain how we make the connection between both.
First, we compute the upper bound on the tpr for a given $(\epsilon,\delta)$-DP privacy budget at a given fpr using Equation A175 reformulated from Kairouz et al. (2015) below:
@@
Next, we use the linear model from Section 4.3 to solve for the minimum $S$ predicted to be required given $C=2$ classes in our example. The coefficients can be found in Table A8 for the average case and Section 4.4 for the worst-case. We solve the tpr from the linear model as
@@
Now, we find the minimum $S$ that the tpr from Equation A177 upper bounds the tpr of Equation A175 as

### v5

@@ B.1 Formulating RMIA
Let $\hat{\mu}_{\bm{a},\bm{b}}$ and $\hat{\sigma}_{\bm{a},\bm{b}}$ denote, respectively, the mean and standard deviation of $t_{\bm{b}}=\ell(\mathcal{M}(\bm{b}),y_{\bm{b}})$ estimated from shadow models, where $\bm{a}$ denotes which of $(\bm{x},y_{\bm{x}})$ and $(\bm{z},y_{\bm{z}})$ is in the training set. By Equation 11 in (Zarifzadeh et al., 2024), the pairwise likelihood ratio is given as
@@
$\mathrm{Score}_{\mathrm{RMIA}}(\bm{x})=\Pr_{(\bm{z},y_{\bm{z}})\sim{\mathbb{D}}}\left(\mathrm{LR}(\bm{x},\bm{z})\geq\gamma\right),$ (A80)
@@ B.2 Per-example RMIA
[…]
@@
When $(\bm{x},y_{\bm{x}})\in\mathcal{D}_{\mathrm{target}}$ and $(\bm{z},y_{\bm{z}})\notin\mathcal{D}_{\mathrm{target}}$, $t_{\bm{x}}=\mu_{\bm{x},\bm{x}}+\sigma_{\bm{x}}Z$ and $t_{\bm{z}}=\mu_{\bm{x},\bm{z}}+\sigma_{\bm{z}}Z$. Thus, $\lambda$ becomes
@@
$\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ $=\Pr_{Z}\left(\Pr_{\bm{z}}(\lambda_{\bm{x}}\geq\log\gamma)\geq\tau\right)$ (A117)
@@
Note that, unlike per-example LiRA, we must assume that the attacker has access to the underlying distribution for the optimal RMIA as the Equations˜A91 and A92 depend on the parameters computed from shadow models.
@@ B.3 RMIA power-law
[…]
@@
To apply Lemma˜8, we will calculate $\mathbb{E}_{\bm{z}}[|q|]$ and $\mathbb{E}_{\bm{z}}[|A|]$. Let $\mathcal{D}_{\mathrm{target}}=\{(\bm{x}_{j,1},j),\dots,(\bm{x}_{j,S},j)\}_{j=1}^{C}$. Let $s_{y_{a}}^{(\bm{b})}(\bm{a})$ denote the score of $(\bm{a},y_{\bm{a}})$ when $\mathcal{D}_{\mathrm{target}}$ contains $(\bm{b},y_{\bm{b}})$ but not the other example. Using similar argument as in the proof of Theorem˜2, we have
@@
[…]
@@
[…]
@@
$=$ $\frac{1}{C}\mathbb{E}_{\bm{z}}\left[\left.\left(\frac{\langle\bm{x},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{x}^{T}\Sigma\bm{x}}}\right)^{2}+\left(\frac{\langle\bm{z},\bm{x}-\bm{z}\rangle}{\sqrt{S}\sqrt{\bm{z}^{T}\Sigma\bm{z}}}\right)^{2}\ \right|\ y_{\bm{z}}=y_{\bm{x}}\right]$ (A142)
@@
Now that from Lemma˜8 we have
@@
We claim that the bound for $\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation˜A159) is as tight as that for $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})$ (Equation˜A156) for sufficiently large $S$. Let us denote
@@
Hence $\textsc{tpr}_{\mathrm{RMIA}}(\bm{x})-\textsc{fpr}_{\mathrm{RMIA}}(\bm{x})=O(1/\sqrt{S})$. On the other hand, from Equations˜A156 and A159 we have for sufficiently large $S$
@@
where $c_{1}$ is some constant. Since for large $S$, $\mathbb{E}_{\bm{z}}[|\hat{\lambda}_{\bm{z}}|]$ and $\tau$ are independent of $S$, Equation˜A173 scales as $O(1/\sqrt{S})$.
@@
As for the LiRA power-law, bounding $\bm{x}-m_{\bm{x}}||$ and $\bm{z}-m_{\bm{z}}||$ will provide a worst-case upper bound for which the power-law holds.
@@ C.3 Datasets
Table˜A2 shows the datasets used in the paper. We base our experiments on a subset of the the few-shot benchmark VTAB (Zhai et al., 2019) that achieves a classification accuracy $80\%$ and thus would considered to be used by a practitioner. Additionally, we add CIFAR10 which is not part of the original VTAB benchmark.
@@ Appendix D Additional results
D.1 Additional results for Section˜4
This Section contains additional results for Section˜4.
@@ D.1.1 Vulnerability as a function of shots
This section displays additional results to Figure˜1 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures˜A.1, A3 and A4.
(a) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
(b) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
(c) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
(d) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
(e) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
(f) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
@@ D.1.2 Vulnerability as a function of the number of classes
This section displays additional results to Figure˜2 for fpr $\in\{0.1,0.01,0.001\}$ for ViT-B and R-50 in in Figures˜A.2, A5 and A6.
(a) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
(b) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.1$
(c) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
(d) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.01$
(e) ViT-B Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
(f) R-50 Head $\textsc{tpr}-\textsc{fpr}$ at $\textsc{fpr}=0.001$
@@ D.1.3 Data for FiLM and from scratch training
Table A7: MIA vulnerability data used in Figure˜5(b). Note that the data from Carlini et al. (2022) is only partially tabular, thus we estimated the tpr at fpr from the plots in the Appendix of their paper.
@@ D.1.4 Predicting dataset vulnerability as function of $S$ and $C$
This section provides additional results for the model based on Equation˜13
Table A8: Results for fitting Equation˜13 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
@@
Figure˜A.3 shows the performance for all considered fpr.
(a) ($(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
Figure A.3: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation˜13 fitted Table˜A3 (ViT-B). The triangles show the median $\textsc{tpr}-\textsc{fpr}$ for the train set (ViT-B; Table˜A3) and circle the test set (R-50; Table˜A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
D.2 Simpler variant of the prediction model
The prediction model in the main text (Equation˜13) avoids predicting $\textsc{tpr}<\textsc{fpr}$ in the tail when $S$ is very large. In this section, we analyse a variation of the regression model that is simpler and predicts $\log_{10}(\textsc{tpr})$ instead of $\log_{10}(\textsc{tpr}-\textsc{fpr})$. This variation fits worse to the empirical data and will predict $\textsc{tpr}<\textsc{fpr}$ for high $S$.
The general form this variant can be found in Equation˜A191, where $\beta_{S},\beta_{C}$ and $\beta_{0}$ are the learnable regression parameters.
Table˜A9 provides tabular results on the performance of the variant.
Table A9: Results for fitting Equation˜A191 with statsmodels Seabold and Perktold (2010) to ViT Head data at fpr $\in\{0.1,0.01,0.001\}$. We utilize an ordinary least squares. The test $R^{2}$ assesses the fit to the data of R-50 Head.
@@ D.2 Simpler variant of the prediction model
Figure˜A.4 plots the performance of the variant similar to Figure˜5(a) in the main text.
@@
Figure A.4: Predicted MIA vulnerability as a function of $S$ (shots) using a model based on Equation˜A191 fitted Table˜A3 (ViT-B). The triangles show the median tpr for the train set (ViT-B; Table˜A3) and circle the test set (R-50; Table˜A4) over six seeds. Note that the triangles and dots for $C=10$ are for EuroSAT.
@@ D.3 Empirical results for RMIA
Figures˜A.5, A.6 and A.7 report additional results for RMIA Zarifzadeh et al. (2024).
(a) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
Figure A.5: RMIA (Zarifzadeh et al., 2024) vulnerability ($\textsc{tpr}-\textsc{fpr}$ at fixed fpr) as a function of $S$ (shots) when attacking a ViT-B Head fine-tuned without DP on different datasets. We observe at power-law relationship but especially at low fpr the relationship is not as clear as with LiRA (compare to Figure˜A.1). The solid line displays the median and the error bars the minimum of the lower bounds and maximum of the upper bounds for the Clopper-Pearson CIs over six seeds.
(a) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
Figure A.6: Predicted MIA vulnerability ($(\textsc{tpr}-\textsc{fpr})$ at fpr) based on LiRA vulnerability data as a function of $S$ (shots) in comparison to observed RMIA (Zarifzadeh et al., 2024) vulnerability on the same settings. The triangles show the highest tpr when attacking (ViT-B Head) with RMIA over six seeds (datasets: Patch Camelyon, EuroSAT and CIFAR100). Especially at $\textsc{fpr}=0.1$ the relationship behaves very similar for both MIAs, but RMIA shows more noisy behavior at lower fpr.
(a) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.1$
(b) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.01$
(c) $(\textsc{tpr}-\textsc{fpr})$ at $\textsc{fpr}=0.001$
Figure A.7: LiRA and RMIA vulnerability ($(\textsc{tpr}-\textsc{fpr})$) as a function of shots ($S$) when attacking a ViT-B Head fine-tuned without DP on different datasets. For better visibility, we split the datasets into two panels. We observe the power-law for both attacks, but the RMIA is more unstable than LiRA. The lines display the median over six seeds.
@@
Table˜A10 displays the tabular results for Figure˜6 in Section˜4.4.
Table A10: Tabular results for Figure˜6 on when attacking a ViT-B (Head) fine-tuned on PatchCamelyon. We display the median over six seeds at $\textsc{fpr}=0.1$.
@@
In Section˜4.5, we compare the results of our empirical models of vulnerability (Sections˜4.3 and 4.4) to DP bounds. Below we explain how we make the connection between both.
First, we compute the upper bound on the tpr for a given $(\epsilon,\delta)$-DP privacy budget at a given fpr using Equation˜A192 reformulated from Kairouz et al. (2015) below:
@@
Next, we use the linear model from Section˜4.3 to solve for the minimum $S$ predicted to be required given $C=2$ classes in our example. The coefficients can be found in Table˜A8 for the average case and Section˜4.4 for the worst-case. We solve the tpr from the linear model as
@@
Now, we find the minimum $S$ that the tpr from Equation˜A194 upper bounds the tpr of Equation˜A192 as